OpenStudy (anonymous):

function f(X) =(1/x) -2 is one to one find the inverse

OpenStudy (anonymous):

y = 1/x - 2 multiply by x: yx = 1 - 2x yx+2x = 1 x(y+2) = 1 x = 1/(y+2) f^-1(x) = 1/(x+2).

OpenStudy (anonymous):

x= 1/y -2 x+2= 1/y y= 1/(x+2)